OpenStudy (anonymous):

WILL AWARD MEDAL! A right triangle has base (x - 7) units and height (2x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle?

OpenStudy (anonymous):

= (x-2)^2 + (2x-10)^2

OpenStudy (anonymous):

use Pythogarass theorem (x-7)^2 +(2x-10)^2

OpenStudy (anonymous):



OpenStudy (anonymous):

\[l^2 = h^2 + w^2\]

OpenStudy (anonymous):

i get the theory i just need to know if (x-7)^2 +(2x-10)^2 is the end result?

OpenStudy (anonymous):

no you have to square them and simplify the results

OpenStudy (anonymous):

i tired but i dont think im getting the right answer. i need help

OpenStudy (anonymous):

it`ll be 5x^2 -54x +194

OpenStudy (anonymous):

Thanks man. I appreciate the help

OpenStudy (anonymous):

some how i got 5x^2-54+149

OpenStudy (anonymous):

*54x

OpenStudy (primeralph):

That's what it should be.

OpenStudy (anonymous):

thanks for the input.